OpenStudy (anonymous):

Could some help me solve this? x2 + 15x = -56 x = 8 and x = 7 x = -8 and x = -7 x = -8 and x = 7 x = 8 and x = -7

OpenStudy (anonymous):

I think it's the first choice, but I'm not sure

OpenStudy (anonymous):

the first choice is correct ^_^

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

oh, sorry sorry, haha its the second choice

OpenStudy (anonymous):

because (x+8)(x+7) then we do 0=(x+8), so x=-8 0=(x+7), so x=-7

OpenStudy (anonymous):

Oh Okay, I get what your saying! Thanks for the help ^-^